### v1

# Rudin-Shapiro function along irreducible polynomials over finite fields

Abstract.
Let $\mathbb{F}_{q}$ be the finite field of $q$ elements. We define the Rudin-Shapiro function $R$ on monic polynomials $f=t^{n}+f_{n-1}t^{n-1}+\dots+f_{0}\in\mathbb{F}_{q}[t]$ over $\mathbb{F}_{q}$ by
$R(f)=\sum_{i=1}^{n-1}f_{i}f_{i-1}.$
We investigate the distribution of the Rudin-Shapiro function along the irreducible polynomials. We show, that the number of irreducible polynomials $f$ with $R(f)=\gamma$ for any $\gamma\in\mathbb{F}_{q}$ is asymptotically $q^{n-1}/n$ as $n\rightarrow\infty$.
Key words and phrases:
irreducible polynomials, finite fields, exponential sums, Rudin-Shapiro function
2020 Mathematics Subject Classification:
11A63,11T23, 11T30
1. Introduction
In recent years, many spectacular results have been obtained on important problems combining some arithmetic properties of the integers and some conditions on their digits in a given basis, see for example [2, 3, 4, 12, 13, 15, 18, 22, 26]. In particular, Mauduit and Rivat [13] showed that the Rudin-Shapiro sequence along prime numbers is well-distributed.
A natural question is to study analog problems in finite fields, see for example [5, 7, 8, 6, 9, 16, 11, 14, 19, 20, 23, 25, 24]. Many of these problems can be solved for finite fields although their analogs for integers are actually out of reach.
In particular, Dartyge and Sárközy [7] investigated the distribution of the Thue-Morse function in finite fields over polynomial values (see also [6]) and Dartyge, Mérai and Winterhof [6] studied the frequency problem of the Rudin-Shapiro function in finite fields over polynomial values.
Moreover, Porritt [21] studied the distribution of linear forms of irreducible polynomials as a generalization of the Thue-Morse function for polynomials.
In this paper, we investigate the frequency problem for the Rudin-Shapiro function along irreducible polynomials.
Let $q$ be a prime power and $\mathbb{F}_{q}$ be the finite field of $q$ elements. For a monic polynomial $f\in\mathbb{F}_{q}[t]$ of degree $n\geq 2$, define the Rudin-Shapiro function $R(f)$ as
(1) $R(f)=\sum_{i=1}^{n-1}f_{i}f_{i-1}\quad\text{with }f=t^{n}+f_{n-1}t^{n-1}+\dots+f_{0}\in\mathbb{F}_{q}[t].$
Let ${\mathcal{P}}(n)$ be the set of monic irreducible polynomials of degree $n$. It is well-known, that
(2) $\frac{q^{n}}{n}-2\frac{q^{n/2}}{n}\leq\#{\mathcal{P}}(n)\leq\frac{q^{n}}{n}$
(see for example [10, Theorem 3.25]). For $\gamma\in\mathbb{F}_{q}$, let
${\mathcal{R}}_{n}(\gamma)=\{f\in{\mathcal{P}}(n):R(f)=\gamma\}.$
Our goal is to prove that the size of $R_{n}(\gamma)$ is asymptotically the same for all $\gamma$.
Our main result is the following theorem.
Theorem 1.
Let $q$ be an odd prime power. There exists a $\delta>0$ such that for any $\gamma\in\mathbb{F}_{q}$ and $n\geq 3$ we have
#$\#{\mathcal{R}}_{n}(\gamma)=\frac{\#{\mathcal{P}}(n)}{q}+O\left(n^{3/2}q^{(1-\delta)n}\right),$
where the implied constant may depend only on $q$. One can choose $\delta=1/30$.
We remark, that for $r=2$ the result follows from the Weil bound, see comment before [17, Problem 8].
2. Auxiliary results
2.1. Notations
For a positive integer $n$, let ${\mathcal{G}}(n)\subset\mathbb{F}_{q}[t]$ be the set of polynomials of degree $n$ and ${\mathcal{M}}(n)\subset{\mathcal{G}}(n)$ be the set of monic polynomials of degree $n$.
Let us identify ${\mathcal{G}}(n)$ with the $n+1$ dimensional linear space $\mathbb{F}_{q}^{n+1}$ via
(3) $f=f_{n}t^{n}+\dots+f_{0}\in{\mathcal{G}}(n)\quad\longleftrightarrow\quad\mathbf{f}=(f_{0},\dots,f_{n})\in\mathbb{F}_{q}^{n+1}.$
Moreover, let us also identify $R$ with the $(n+1)\times(n+1)$ matrix $R=(R_{i,j})$ defined by
(4) $R_{i,j}=\left\{\begin{array}[]{cl}1&\text{if $i=j+1$}\\ 0&\text{otherwise}.\end{array}\right.$
Then we have $R(f)=\mathbf{f}^{T}R\mathbf{f}$, where $R(f)$ is defined as (1).
For given functions $F$ and $G$, the notations $F\ll G$, $G\gg F$ and $F=O(G)$ are all equivalent to the statement that the inequality $|F|\leq c|G|$ holds with some constant $c>0$. Throughout the paper, any implied constants in symbols $O$, $\ll$ and $\gg$ may depend only on $q$.
2.2. Quadratic forms of polynomials
Let $Q$ be a quadratic form on $\mathbb{F}_{q}^{n}$. Then, $Q$ can be written as $Q(x)=x^{T}Mx$ for some matrix $M\in\mathbb{F}_{q}^{n\times n}$. The rank of $Q$ is the rank of the matrix $M$. The first result we need is the following bound. For completeness, we also include the proof.
Lemma 2.
Let $q$ be and odd prime power and let $Q$ be a quadratic and $L$ be a linear form on $\mathbb{F}_{q}^{n}$. Assume, that $Q$ has rank at least $r$. Then
$\left|\sum_{x\in\mathbb{F}_{q}^{n}}\psi(Q(x)+L(x))\right|\leq q^{n-r/2}$
for any non-trivial additive character $\psi$ of $\mathbb{F}_{q}$.
Proof.
Write $Q(x)=x^{T}Mx$ for some symmetric matrix $M\in\mathbb{F}_{q}^{n}$ of rank at least $r$ and let $B$ be the corresponding bilinear form
$B(x,y)=x^{T}My.$
Write also $B_{x}(y)=B(x,y)$. Let $K\leq\mathbb{F}_{q}^{n}$ of vectors $x$ such that $B_{x}=0$. It’s codimension is the rank of $M$ and thus of the quadratic form $Q$.
We have
$\left|\sum_{x\in\mathbb{F}_{q}^{n}}\psi(Q(x)+L(x))\right|^{2}$ $=\sum_{x,h\in\mathbb{F}_{q}^{n}}\psi(Q(x+h)+L(x+h))\overline{\psi(Q(x)+L(x))}$
$=\sum_{h\in\mathbb{F}_{q}^{n}}\psi(Q(h)+L(h))\sum_{x\in\mathbb{F}_{q}^{n}}\psi(2B_{h}(x)).$
If $h\not\in K$, then $B_{h}$ is a non-zero linear form and thus
$\sum_{x\in\mathbb{F}_{q}^{n}}\psi(2B_{h}(x))=\sum_{x\in\mathbb{F}_{q}^{n}}\psi(x)=0.$
For $h\in K$, we have
$\sum_{x\in\mathbb{F}_{q}^{n}}\psi(2B_{h}(x))=q^{n}.$
The assertion follows by noticing #$\#K=q^{n-r}$. ∎
Lemma 3.
Let $k<n/2$ and $a,b\in{\mathcal{M}}(k)$. For any fixed $b\in{\mathcal{M}}(k)$, let $r_{b}$ be the number of $a\in{\mathcal{M}}(k)$ such that the map $h\mapsto R(ah)-R(bh)$ with $h\in{\mathcal{M}}(n-k)$ has rank at most $n-2k$. Then $r_{b}=O(q^{(k+1)/2})$, where the implied constant is absolute.
Proof.
For a given polynomial $f\in{\mathcal{G}}(k)$, write $[f]\in\mathbb{F}_{q}^{(n+1)\times(n-k+1)}$ be the matrix
$[f]_{i,j}=\left\{\begin{array}[]{cl}f_{i-j}&\text{if $i\leq j$}\\ 0&\text{otherwise}.\end{array}\right.$
Observe, that if $g\in{\mathcal{G}}(n-k)$, then writing $\mathbf{g}=(g_{0},\dots,g_{n-k})^{T}$, $[f]\mathbf{g}$ corresponds to the product polynomial $f\cdot g$ via (3). It yields $R(fg)=\mathbf{g}^{T}[f]^{T}R[f]\mathbf{g}$, where the matrix $R$ is defined as (4). Then the map $h\mapsto R(ah)-R(bh)$ corresponds to $M_{a,b}=[a]^{T}R[a]-[b]^{T}R[b]$.
First write $A=[a]^{T}R[a]\in\mathbb{F}_{q}^{(n-k+1)\times(n-k+1)}$. Observe, that the entries of $A$ at position $(i,j)$ with $i\leq j\leq n-k$ are
$A_{i,j}=\sum_{\ell=0}^{k-j+i-1}a_{\ell}a_{\ell+j-i+1}.$
Specially, $A_{i,j}=0$ for $j>i+k$.
Define $S=(s_{i,j})\in\mathbb{F}_{q}^{(k+1)\times(k+1)}$ as
$s_{i,j}=\left\{\begin{array}[]{cl}1&\text{if $i=j+1$}\\ 0&\text{otherwise}.\end{array}\right.$
Then
$A_{i,j}=\mathbf{a}^{T}S^{j-i+1}\mathbf{a},\quad\text{where }\mathbf{a}=(a_{0},\dots,a_{k})^{T}\in\mathbb{F}_{q}^{k+1}.$
Similarly, writing $b=[b]^{T}R[b]\in\mathbb{F}_{q}^{(n-k+1)\times(n-k+1)}$, then
$B_{i,j}=\mathbf{b}^{T}S^{j-i+1}\mathbf{b},\quad\text{where }\mathbf{b}=(a_{0},\dots,b_{k})^{T}\in\mathbb{F}_{q}^{k+1}.$
It yields,
$(M_{a,b})_{i,i+d}=\left\{\begin{array}[]{cc}\mathbf{a}^{T}S^{d+1}\mathbf{a}-\mathbf{b}^{T}S^{d+1}\mathbf{b}&\text{if $0\leq d<k$}\\ 0&\text{if $d\geq k$}.\end{array}\right.$
Assume, that $\mathbf{a}$ is so that
(5) $\mathbf{a}^{T}S^{d+1}\mathbf{a}-\mathbf{b}^{T}S^{d+1}\mathbf{b}\neq 0$
for some $0\leq d<k$, and let $d^{*}$ be the maximal index such that (5) holds with $d^{*}$ instead if $d$. Then the submatrix of $(M_{a,b})$ formed by $(M_{a,b})_{i,j}$ with $i=1,\dots,n-2k+1$ and $j=d^{*},\dots,d^{*}+n-2k+1$ is a lower-triangular matrix with non-zero elements in the diagonal, thus it has rank $n-2k+1$.
Thus, to prove the result, it is enough to estimate the number $N$ of $a\in{\mathcal{M}}(k)$ such that (5) does not hold for any $0\leq d<k$, that is
$\mathbf{a}^{T}S^{d+1}\mathbf{a}=\mathbf{b}^{T}S^{d+1}\mathbf{b}\quad 0\leq d<k.$
Let $\psi$ be a non-trivial additive character of $\mathbb{F}_{q}$. Then
$N$ $=\frac{1}{q^{k}}\sum_{a\in{\mathcal{M}}(k)}\sum_{t_{1},\dots,t_{k}}\psi\left(\sum_{i=1}^{k}t_{i}\left(\mathbf{a}^{T}S^{i}\mathbf{a}-\mathbf{b}^{T}S^{i}\mathbf{b}\right)\right)$
$=\frac{1}{q^{k}}\sum_{t_{1},\dots,t_{k}}\psi\left(-\mathbf{b}^{T}\left(\sum_{i=1}^{t}t_{i}S^{i}\right)\mathbf{b}\right)\sum_{a\in{\mathcal{M}}(k)}\psi\left(\mathbf{a}^{T}\left(\sum_{i=1}^{k}t_{i}S^{i}\right)\mathbf{a}\right).$
The entries of
(6) $t_{1}S+t_{2}S^{2}+\dots+t_{k}S^{k}\in\mathbb{F}_{q}^{(k+1)\times(k+1)}$
at position $(i,i+u)$ is $t_{u}$ for $1\leq u\leq k$ and 00 for $u>k$. Thus the rank of (6) is $k+1-\min\{1\leq u\leq k:t_{u}\neq 0\}$.
As the rank of (6) is reduced at most one if we consider it on the affine space of monic polynomials $a$, we get by Lemma 2,
$N$ $\leq 1+\frac{1}{q^{k}}\sum_{u=1}^{k}\sum_{\begin{subarray}{c}t_{v}\in\mathbb{F}_{q}\\ u<v\leq k\end{subarray}}\sum_{t_{u}\in\mathbb{F}_{q}^{*}}\left|\sum_{a\in{\mathcal{M}}(k)}\psi\left(\mathbf{a}^{T}\left(\sum_{i=u}^{k}t_{i}S^{i}\right)\mathbf{a}\right)\right|$
$\ll\frac{1}{q^{k}}\sum_{u=1}^{k}q^{k-u+1}q^{k-(k-u)/2}\ll q^{(k+1)/2}.$
Lemma 4.
Let $a\in{\mathcal{M}}(k)$. Then for $h\in{\mathcal{M}}(n-k)$, the bilinear map $h\mapsto R(ah)$ has rank at least $n-k-1$.
Proof.
Write $a=t^{u}b$ with $t\nmid b$ and $b=t^{k-u}+b_{k-u-1}t^{k-u-1}+\dots+b_{0}$. Moreover, define $B=(B_{i,j})\in\mathbb{F}_{q}^{(n-u)\times(n-k)}$ as
$B_{i,j}=\left\{\begin{array}[]{cl}b_{i-j}&\text{if $i\geq j$}\\ 0&\text{otherwise},\end{array}\right.$
and $S=(S_{i,j})\in\mathbb{F}_{q}^{(n-u)\times(n-u)}$ as
$S_{i,j}=\left\{\begin{array}[]{cl}1&\text{if $i=j+1$}\\ 0&\text{otherwise}.\end{array}\right.$
Then, the bilinear map $R(ah)$ corresponds to the matrix $B^{T}SB$.
Write
$B=\left(\begin{matrix}B^{\prime}\\ P\end{matrix}\right)\quad\text{and}\quad S=\left(\begin{matrix}S_{1}&\mathbf{0}\\ S_{2}&S_{3}\end{matrix}\right)$
with $B^{\prime}\in\mathbb{F}_{q}^{(n-k)\times(n-k)}$, $P\in\mathbb{F}_{q}^{(k-u)\times(n-k)}$, $S_{1}\in\mathbb{F}_{q}^{(n-k)\times(n-k)}$, $S_{2}\in\mathbb{F}_{q}^{(k-u)\times(n-k)}$ and $S_{3}\in\mathbb{F}_{q}^{(k-u)\times(k-u)}$.
As $t\nmid b$, $b_{0}\neq 0$, so $B^{\prime}$ is a lower triangular matrix with non-zero element in the diagonal, thus has full rank $n-k$. Also, $S_{1}$ has rank $n-k-1$. Thus
$B^{T}SB=\left(\begin{matrix}(B^{\prime})^{T}S_{1}B^{\prime}&\mathbf{0}\\ P^{T}S_{2}B^{\prime}&P^{T}S_{3}P\end{matrix}\right)$
has the submatrix $(B^{\prime})^{T}S_{1}B^{\prime}$ of rank $n-k-1$. ∎
2.3. Divisor bounds
Let $\tau(f)$ denote the number of monic divisors of $f\in\mathbb{F}_{q}[t]$. We need the following bound on the second moment of $\tau$, see [1, Lemma 8].
Lemma 5.
We have
$\sum_{f\in{\mathcal{M}}(n)}\tau(f)^{2}\leq 4n^{3}q^{n}.$
2.4. Vaughan’s identity for function fields
In this section we revise Vaughan’s identity for function fields. To do that, let us define von Mangoldt function $\Lambda$ in $\mathbb{F}_{q}[t]$ by
$\Lambda(f)=\left\{\begin{array}[]{cl}\deg P&\text{if $f=P^{r}$ for some irreducible polynomial $P$;}\\ 0&\text{otherwise.}\end{array}\right.$
Then we have the following result.
Lemma 6.
Let $\Psi:\mathbb{F}_{q}[t]\rightarrow\mathbb{C}$ with $|\Psi(f)|\leq 1$. If $1\leq u,v\leq n$ two integers with $u+v<n$, then
$\sum_{f\in{\mathcal{M}}(n)}\Lambda(f)\Psi(f)\ll n\Sigma_{1}+n^{5/2}q^{n-(u+v)/2}\Sigma_{2}^{1/2}$
with
$\Sigma_{1}$ $=\sum_{|g|\leq q^{u+v}}\left|\sum_{|h|=q^{n}/|g|}\Psi(gh)\right|$
$\Sigma_{2}$ $=\max_{v\leq i\leq n-u}\max_{|g_{1}|=q^{n-i}}\sum_{|g_{2}|=q^{n-i}}\left|\sum_{|h|=q^{i}}\Psi(hg_{1})\overline{\Psi(hg_{2})}\right|,$
where the summations are taken over monic polynomials and the implied constant is absolute.
Proof.
Let $\mu$ be the Möbius function on $\mathbb{F}_{q}[t]$ defined by
$\mu(f)=\left\{\begin{array}[]{cl}(-1)^{k}&\text{if $f=P_{1}\dots P_{k}$ for some distinct }\\ &\text{\quad irreducible polynomials $P_{1},\dots,P_{k}$;}\\ 0&\text{otherwise.}\end{array}\right.$
Then, one can write
$\Lambda(f)=\sum_{a\mid f}\mu(a)\deg(f/a),$
where the sum is taken over only monic polynomials. Then
$\Lambda(f)=\sum_{\begin{subarray}{c}a\mid f\\ |a|\leq q^{u}\end{subarray}}\mu(a)\deg(f/a)-\underset{\begin{subarray}{c}ab\mid f\\ |a|\leq q^{u},|b|\leq q^{v}\end{subarray}}{\sum\ \sum}\mu(b)\Lambda(b)+\underset{\begin{subarray}{c}ab\mid f\\ |a|>q^{u},|b|>q^{v}\end{subarray}}{\sum\ \sum}\mu(a)\Lambda(b).$
It yields
$\sum_{f\in{\mathcal{M}}(n)}\Lambda(f)\Psi(f)=S_{1}-S_{2}+S_{3},$
where
$S_{1}$ $=\underset{\begin{subarray}{c}|ab|=q^{n}\\ |a|\leq q^{u}\end{subarray}}{\sum\ \sum}\mu(a)\deg(b)\Psi(ab),$
$S_{2}$ $=\underset{\begin{subarray}{c}|abc|=q^{n}\\ |a|\leq q^{u},|b|\leq q^{v}\end{subarray}}{\sum\ \sum\ \sum}\mu(a)\Lambda(b)\Psi(abc),$
$S_{3}$ $=\underset{\begin{subarray}{c}|abc|=q^{n}\\ |a|>q^{u},|b|>q^{v}\end{subarray}}{\sum\ \sum\sum}\mu(a)\Lambda(b)\Psi(abc).$
For $S_{1}$, we have
$S_{1}=\sum_{|a|\leq q^{u}}\mu(a)(n-\deg(a))\sum_{\begin{subarray}{c}|b|=q^{n}/|a|\end{subarray}}\Psi(ab)\ll n\sum_{|a|\leq q^{u}}\left|\sum_{\begin{subarray}{c}|b|=q^{n}/|b|\end{subarray}}\Psi(ab)\right|.$
For $S_{2}$, observe, that
$\sum_{b\mid z}\Lambda(b)=\deg(z).$
Then, we have
$\begin{split}S_{2}&=\sum_{r=n-u-v}^{n}\left(\underset{\begin{subarray}{c}|a|\leq q^{u}\,|b|\leq q^{v}\\ |ab|=q^{n-r}\end{subarray}}{\sum\ \sum}\mu(a)\Lambda(b)\right)\sum_{|c|=q^{r}}\Psi(abc)\\ &\ll(u+v)\sum_{|z|\leq q^{u+v}}\left|\sum_{|c|=q^{n}/|z|}\Psi(zc)\right|.\end{split}$
For $S_{3}$, observe, that
$\sum_{a\mid c}\mu(a)=\left\{\begin{array}[]{cl}1&\text{if $c=1$};\\ 0&\text{otherwise}.\end{array}\right.$
Then, we have
$S_{3}=\sum_{q^{v}\leq|b|\leq q^{n-u}}\Lambda(b)\sum_{|c|=q^{n}/|b|}\left(\sum_{\begin{subarray}{c}a\mid c\\ |a|>q^{u}\end{subarray}}\mu(a)\right)\Psi(bc).$
For $c\in\mathbb{F}_{q}[t]$, write
$\beta_{c}=\sum_{\begin{subarray}{c}a\mid c\\ |a|>q^{u}\end{subarray}}\mu(a).$
Then, by the Cauchy-Schwarz inequality,
$\left|S_{3}\right|^{2}\leq\left(\sum_{q^{v}\leq|b|\leq q^{n-u}}\Lambda(b)^{2}\right)\sum_{q^{v}\leq|b|\leq q^{n-u}}\left|\sum_{|c|=q^{n}/|b|}\beta_{c}\Psi(bc)\right|^{2}.$
For the first term, we have
$\sum_{q^{v}\leq|b|\leq q^{n-u}}\Lambda(b)^{2}\leq n\sum_{|b|\leq q^{n-u}}\Lambda(b)\leq nq^{n-u}.$
For the second term,
$\sum_{q^{v}\leq|b|\leq q^{n-u}}\left|\sum_{|c|=q^{n}/|b|}\beta_{c}\Psi(bc)\right|^{2}$
$=\sum_{v\leq i\leq n-u}\sum_{|c_{1}|=q^{n-i}}\sum_{|c_{2}|=q^{n-i}}\beta_{c_{1}}\overline{\beta_{c_{2}}}\sum_{|b|=q^{i}}\Psi(bc_{1})\overline{\Psi(bc_{2})}$
$\ll\sum_{v\leq i\leq n-u}\sum_{|c_{1}|=q^{n-i}}|\beta_{c_{1}}|^{2}\sum_{|c_{2}|=q^{n-i}}\left|\sum_{|b|=q^{i}}\Psi(bc_{1})\overline{\Psi(bc_{2})}\right|$
$\ll\sum_{v\leq i\leq n-u}\left(\sum_{|c_{1}|=q^{n-i}}|\beta_{c_{1}}|^{2}\right)\max_{|c_{1}|=q^{n-i}}\sum_{|c_{2}|=q^{n-i}}\left|\sum_{|b|=q^{i}}\Psi(bc_{1})\overline{\Psi(bc_{2})}\right|$
by $|\beta_{c_{1}}\overline{\beta_{c_{2}}}|\leq|\beta_{c_{1}}|^{2}/2+|\beta_{c_{2}}|^{2}/2$ and by the symmetry.
By Lemma 5,
$\sum_{|c_{1}|=q^{n-i}}|\beta_{c_{1}}|^{2}\ll\sum_{|c_{1}|=q^{n-i}}|\tau(c_{1})|^{2}\ll n^{3}q^{n-i}.$
Thus
$|S_{3}|^{2}\ll n^{5}q^{2n-u-v}\max_{v\leq i\leq n-u}\max_{|c_{1}|=q^{n-i}}\sum_{|c_{2}|=q^{n-i}}\left|\sum_{|b|=q^{i}}\Psi(bc_{1})\overline{\Psi(bc_{2})}\right|.$
3. Proof of Theorem 1
Let $\gamma\in\mathbb{F}_{q}$. By the orthogonality of additive characters, we have
(7) ${\mathcal{R}}(\gamma)=\frac{1}{q}\sum_{\psi}\sum_{f\in{\mathcal{P}}(n)}\psi(R(f)-\gamma),$
where the first summation is taken over all additive characters of $\mathbb{F}_{q}$. For the trivial additive character $\psi_{0}$, we have
(8) $\frac{1}{q}\sum_{f\in{\mathcal{P}}(n)}\psi_{0}(R(f)-\gamma)=\frac{\#{\mathcal{P}}(n)}{q}$
which gives the main term.
The number of irreducible polynomials $f$ of degree $\deg f<n$ with $\deg f\mid n$ is $O(q^{n/2})$ by (2), thus for non-trivial additive characters $\psi\neq\psi_{0}$, we have
(9) $\begin{split}\sum_{f\in{\mathcal{P}}(n)}\psi(R(f))&=\frac{1}{n}\sum_{f\in{\mathcal{P}}(n)}\Lambda(f)\psi(R(f))\\ &=\frac{1}{n}\sum_{f\in{\mathcal{M}}(n)}\Lambda(f)\psi(R(f))+O(q^{n/2}).\end{split}$
Put
(10) $u=\left\lceil 3n/10\right\rceil\quad\text{and}\quad v=\left\lceil 19n/30\right\rceil.$
Then, by Lemma 6,
(11) $\sum_{f\in{\mathcal{M}}(n)}\Lambda(f)\psi(R(f))\ll n\Sigma_{1}+n^{5/2}q^{n-(u+v)/2}\Sigma_{2}^{1/2},$
where
$\Sigma_{1}=\sum_{|g|\leq q^{u+v}}\left|\sum_{|h|=q^{n}/|g|}\psi(R(gh))\right|$
and
$\Sigma_{2}=\max_{v\leq i\leq n-u}\max_{|g_{1}|=q^{n-i}}\sum_{|g_{2}|=q^{n-i}}\left|\sum_{|h|=q^{i}}\psi(R(hg_{1}))\overline{\psi(R(hg_{2}))}\right|.$
In order to estimate $\Sigma_{1}$, observe, that by Lemma 4, the quadratic polynomial $h\mapsto R(gh)$ has rank at least $n-\deg g-1$. Then, by Lemma 2,
(12) $\Sigma_{1}$ $\ll\sum_{|g|\leq q^{u+v}}q^{n-\deg g-(n-\deg g)/2}$
$=\sum_{|g|\leq q^{u+v}}q^{(n-\deg g)/2}$
$\ll q^{n/2}\sum_{i\leq u+v}q^{i/2}$
$\ll q^{(n+u+v)/2}\ll q^{29n/30}$
using (10).
In order to estimate $\Sigma_{2}$, let us fix $v\leq i\leq n-u$ and consider
$S(g_{1},g_{1})=\sum_{|h|=q^{i}}\psi(R(g_{1}h)-R(g_{2}h)).$
with monic polynomials $g_{1},g_{2}$ of degree $n-i$. For any $g_{2}$, the number of $g_{1}$ such that the bilinear form $h\mapsto R(g_{1}h)-R(g_{2}h)$ has rank at most $n-2\deg g_{1}=2i-n$ is $O(q^{(n-i)/2})$ by Lemma 3. For those polynomials $g_{1}$, we estimate $S(g_{1},g_{2})$ trivially by $q^{i}$. Otherwise, we apply Lemma 2 to get
$S(g_{1},g_{2})\ll q^{i-(2i-n)/2}=q^{n/2}.$
Then
$\Sigma_{2}\ll\max_{v\leq i\leq n-u}\left(q^{(n-i)/2}q^{i}+q^{n-i}q^{n/2}\right)\ll q^{n-u/2}+q^{3n/2-v}.$
It yields
(13) $n^{5/2}q^{n-(u+v)/2}\Sigma_{2}^{1/2}$ $\ll n^{5/2}\left(q^{3n/2-3u/4-v/2}+q^{7n/4-u/2-v}\right)$
$\ll n^{5/2}q^{29n/30}$
using (10). Then, it follows from (9), (11), (12) and (13) that
$\sum_{f\in{\mathcal{P}}(n)}\psi(R(f))\ll n^{3/2}q^{29n/30}.$
Then, the result follows from (7) and (8).
Acknowledgment
The author wishes to thank Arne Winterhof for his valuable comments.
The author was was partially supported by NRDI (National Research Development and Innovation Office, Hungary) grant FK 142960 and by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.