### v2

# Rudin-Shapiro function along irreducible polynomials over finite fields

Abstract.
Let $q$ be an odd prime power and $\mathbb{F}_{q}$ be the finite field of $q$ elements. We define the Rudin-Shapiro function $R$ on monic polynomials $f=t^{n}+f_{n-1}t^{n-1}+\dots+f_{0}\in\mathbb{F}_{q}[t]$ over $\mathbb{F}_{q}$ by
$R(f)=\sum_{i=1}^{n-1}f_{i}f_{i-1}.$
We investigate the distribution of the Rudin-Shapiro function along irreducible polynomials. We show that the number of irreducible polynomials $f$ with $R(f)=\gamma$ for any $\gamma\in\mathbb{F}_{q}$ is asymptotically $q^{n-1}/n$ as $n\rightarrow\infty$.
Key words and phrases:
irreducible polynomials, finite fields, exponential sums, Rudin-Shapiro function
2020 Mathematics Subject Classification:
11A63,11T23, 11T30
1. Introduction
In recent years, many spectacular results have been obtained on important problems combining some arithmetic properties of integers and some conditions on their digits in a given basis, see for example [2, 3, 4, 12, 13, 15, 18, 22, 26]. In particular, Mauduit and Rivat [13] showed that the Rudin-Shapiro sequence along prime numbers is well-distributed.
A natural question is to study analog problems in finite fields, see for example [5, 7, 8, 6, 9, 16, 11, 14, 19, 20, 23, 25, 24]. Many of these problems can be solved for finite fields although their analogs for integers are actually out of reach.
In particular, Dartyge and Sárközy [7] investigated the distribution of the Thue-Morse function in finite fields over polynomial values (see also [6]) and Dartyge, Mérai and Winterhof [6] studied the frequency problem of the Rudin-Shapiro function in finite fields over polynomial values.
Moreover, Porritt [21] studied the distribution of linear forms of irreducible polynomials as a generalization of the Thue-Morse function for polynomials.
In this paper, we investigate the frequency problem for the Rudin-Shapiro function along irreducible polynomials.
Let $q$ be an odd prime power and $\mathbb{F}_{q}$ be the finite field of $q$ elements. For a monic polynomial $f\in\mathbb{F}_{q}[t]$ of degree $n\geq 2$, define the Rudin-Shapiro function $R(f)$ as
$R(f)=\sum_{i=1}^{n-1}f_{i}f_{i-1}\quad\text{with }f=t^{n}+f_{n-1}t^{n-1}+\dots+f_{0}\in\mathbb{F}_{q}[t].$
Let ${\mathcal{P}}(n)$ be the set of monic irreducible polynomials of degree $n$. It is well-known, that
(1) $\frac{q^{n}}{n}-2\frac{q^{n/2}}{n}\leq\#{\mathcal{P}}(n)\leq\frac{q^{n}}{n}$
(see for example [10, Theorem 3.25]). For $\gamma\in\mathbb{F}_{q}$, let
${\mathcal{R}}_{n}(\gamma)=\{f\in{\mathcal{P}}(n):R(f)=\gamma\}.$
Our goal is to prove that the size of ${\mathcal{R}}_{n}(\gamma)$ is asymptotically the same for all $\gamma$.
Our main result is the following theorem.
Theorem 1.
Let $q$ be an odd prime power. There exists a $\delta>0$ such that for any $\gamma\in\mathbb{F}_{q}$ and $n\geq 3$ we have
${\mathcal{R}}_{n}(\gamma)=\frac{\#{\mathcal{P}}(n)}{q}+O\left(n^{3/2}q^{(1-\delta)n+1}\right),$
where the implied constant is absolute. One can choose $\delta=1/28$.
We remark, that for $n=2$ the result follows from the Weil bound, see comment before [17, Problem 8].
By Theorem 1, we have that
${\mathcal{R}}_{n}(\gamma)\sim\frac{\#{\mathcal{P}}(n)}{q}$
if either $n\rightarrow\infty$ or $q\rightarrow\infty$ provided $n>2/\delta$.
2. Outline of the proof
For a positive integer $n$, let ${\mathcal{G}}(n)\subset\mathbb{F}_{q}[t]$ be the set of polynomials of degree $n$ and ${\mathcal{M}}(n)\subset{\mathcal{G}}(n)$ be the set of monic polynomials of degree $n$.
For $\ell\geq 0$, write
$S^{(\ell)}(f)=\sum_{i=\ell}^{n}f_{i}f_{i-\ell}\quad\text{with }f=f_{n}t^{n}+f_{n-1}t^{n-1}+\dots+f_{0}\in\mathbb{F}_{q}[t].$
For $\ell=1$, we simply write
$S(f)=S^{(1)}(f).$
Clearly, for monic polynomials $f$ we have
(2) $R(f)=S(f)-f_{n-1}.$
In order to prove Theorem 1, we introduce Vaughan’s identity for polynomials in Section 4.4. Then the main tool is to investigate single and double exponential sums in terms of $S(f)$ in Section 4.1. The bounds of the exponential sum estimates are given in terms of the rank of quadratic forms which is studied in Sections 4.2 and 4.3. Then we conclude the proof of Theorem 1 in Section 5.
For given functions $F$ and $G$, the notations $F\ll G$, $G\gg F$ and $F=O(G)$ are all equivalent to the statement that the inequality $|F|\leq c|G|$ holds with some constant $c>0$.
3. preliminaries
Let $\tau(f)$ denote the number of monic divisors of $f\in\mathbb{F}_{q}[t]$. First, we state the following bound on $\tau(f)$.
Lemma 2.
For any $\varepsilon>0$ we have
$\tau(f)\ll\min\left\{|f|^{(2+\varepsilon)/\log\deg f},2^{\deg f}\right\},$
where the implied constant can depend on $\varepsilon$.
Proof.
The first bound follows from [16, Lemma 16], while the second bound follows from the fact that a polynomial $f$ has at most $\deg f$ irreducible divisor. ∎
We also have the following bound on the second moment of $\tau$, see [1, Lemma 8].
Lemma 3.
We have
$\sum_{f\in{\mathcal{M}}(n)}\tau(f)^{2}\leq 4n^{3}q^{n}.$
4. Auxiliary results
4.1. Exponential sums of quadratic forms of polynomials
Let $Q$ be a quadratic form on $\mathbb{F}_{q}^{n}$. Then, $Q$ can be written as $Q({\mathbf{x}})={\mathbf{x}}^{T}M{\mathbf{x}}$ for some symmetric matrix $M\in\mathbb{F}_{q}^{n\times n}$. The rank of $Q$ is the rank of the matrix $M$. The first result we need is the following bound. For completeness, we also include the proof.
Lemma 4.
Let $q$ be an odd prime power and let $Q$ be a quadratic form, and let $L$ be a linear form on $\mathbb{F}_{q}^{n}$. Assume that $Q$ has rank at least $r$. Then
$\left|\sum_{{\mathbf{x}}\in\mathbb{F}_{q}^{n}}\psi(Q({\mathbf{x}})+L({\mathbf{x}}))\right|\leq q^{n-r/2}$
for any non-trivial additive character $\psi$ of $\mathbb{F}_{q}$.
Proof.
Write $Q({\mathbf{x}})={\mathbf{x}}^{T}M{\mathbf{x}}$ for some symmetric matrix $M\in\mathbb{F}_{q}^{n\times n}$ of rank at least $r$ and let $B$ be the corresponding bilinear form
$B({\mathbf{x}},{\mathbf{y}})={\mathbf{x}}^{T}M{\mathbf{y}}.$
Write also $B_{\mathbf{x}}({\mathbf{y}})=B({\mathbf{x}},{\mathbf{y}})$. Let $K\leq\mathbb{F}_{q}^{n}$ the subspace of vectors $\mathbf{x}$} such that $B_{\mathbf{x}}=0$. It’s codimension is the rank of $M$ and thus of the quadratic form $Q$.
We have
$\left|\sum_{{\mathbf{x}}\in\mathbb{F}_{q}^{n}}\psi(Q({\mathbf{x}})+L({\mathbf{x}}))\right|^{2}$
$=\sum_{{\mathbf{x}},{\mathbf{h}}\in\mathbb{F}_{q}^{n}}\psi(Q({\mathbf{x}}+{\mathbf{h}})+L({\mathbf{x}}+{\mathbf{h}}))\overline{\psi(Q({\mathbf{x}})+L({\mathbf{x}}))}$
$=\sum_{{\mathbf{h}}\in\mathbb{F}_{q}^{n}}\psi(Q({\mathbf{h}})+L({\mathbf{h}}))\sum_{{\mathbf{x}}\in\mathbb{F}_{q}^{n}}\psi(2B_{\mathbf{h}}({\mathbf{x}})).$
If ${\mathbf{h}}\not\in K$, then $B_{\mathbf{h}}$ is a non-zero linear form and thus
$\sum_{{\mathbf{x}}\in\mathbb{F}_{q}^{n}}\psi(2B_{h}({\mathbf{x}}))=\sum_{{\mathbf{x}}\in\mathbb{F}_{q}^{n}}\psi({\mathbf{x}})=0.$
For ${\mathbf{h}}\in K$, we have
$\sum_{{\mathbf{x}}\in\mathbb{F}_{q}^{n}}\psi(2B_{\mathbf{h}}({\mathbf{x}}))=q^{n}.$
The assertion follows by noticing $K=q^{n-r}$. ∎
4.2. Reversed polynomials
For a polynomial $a\in\mathbb{F}_{q}[t]$ of form $a=a_{n}t^{n}+\dots+a_{1}t+a_{0}$ let $a^{*}$ be the polynomial with reversed coefficients, that is,
$a^{*}(t)=t^{n}a(t^{-1})=a_{0}t^{n}+\dots+a_{n-1}t+a_{n}\in\mathbb{F}_{q}[t].$
First we show, that the quantities $S^{(\ell)}(a)$ can be determined by using the reversed polynomial $a^{*}$ of $a$.
Lemma 5.
For any polynomial $a\in{\mathcal{G}}(n)$, we have
$a^{*}\cdot a=\sum_{\ell=0}^{n}S^{(\ell)}(a)\left(t^{n+\ell}+t^{n-\ell}\right).$
Proof.
The coefficient of the term $t^{n-\ell}$ in $a^{*}\cdot a$ is
$\sum_{i+j=n-\ell}a_{n-i}a_{j}=\sum_{j=0}^{n-\ell}a_{j+\ell}a_{j}=S^{(\ell)}(a).$
Moreover, the coefficient of $t^{n-\ell}$ is the same as the coefficient of $t^{n+\ell}$. ∎
Next, we investigate the following type of quadratic equations in terms of the polynomials $a$ and its reversed $a^{*}$.
Lemma 6.
Let $\varepsilon>0$. For given integer $n$, and polynomial $f\in{\mathcal{G}}(2n)$, let $N(f)$ be the number of polynomials $a\in{\mathcal{G}}(n)$ with
(3) $a^{*}\cdot a=f.$
Then
$N(f)\ll\min\left\{2^{n},q^{(2+\varepsilon)\frac{n}{\log_{q}n}}\right\},$
where the implied constant may depend on $\varepsilon$.
Proof.
Clearly, for any polynomials $a,b\in\mathbb{F}_{q}[t]$, we have
(4) $(a\cdot b)^{*}=a^{*}\cdot b^{*}.$
Now assume that $f$ has the form
$f=b^{*}\cdot b\quad\text{for some }b\in{\mathcal{G}}(n),$
since otherwise the bound trivially holds.
If $a$ is a polynomial satisfying (3), put
$d=\gcd(a,b).$
Then by (4)
$f=d\cdot\frac{b}{d}\cdot d^{*}\cdot\left(\frac{b}{d}\right)^{*},$
and as
$\gcd(a/d,b/d)=1\quad\text{and}\quad\gcd((a/d)^{*},(b/d)^{*})=1$
we must have
$a=d\cdot\left(\frac{b}{d}\right)^{*}.$
That is, the divisor $d$ uniquely determines $a$. Whence, the number of polynomials $a$ with (3) is at most $\tau(b)$. Then the bound follows from Lemma 2. ∎
4.3. Rank of quadratic forms
For a fixed $a\in{\mathcal{M}}(k)$ the map $h\mapsto R(ah)$ with $h\in{\mathcal{G}}(n-k)$ is a quadratic form in the coefficients of $h$. First we investigate its rank.
Lemma 7.
Let $a\in{\mathcal{M}}(k)$ and for $h\in{\mathcal{G}}(n-k)$ consider the map $Q_{a}:h\mapsto S(ah)$, which is a quadratic form in the coefficients of $h$. If $k<n/2$, then it has rank at least $\mathop{\mathrm{rank}}Q_{a}\geq n-k-1$.
Proof.
Let $B$ be the symmetric bilinear form so that $S(h)=B(h,h)$. Specially,
(5) $B(t^{i},t^{j})=\left\{\begin{array}[]{cl}1/2&\text{if }|i-j|=1,\\ 0&\text{otherwise},\end{array}\right.\quad 0\leq i,j\leq n.$
Put
$K=\{g\in{\mathcal{G}}(n-k):h\mapsto B(ag,ah)\text{ is the zero map}\}$
which is a linear subspace of ${\mathcal{G}}(n-k)$. Then
(6) $\mathop{\mathrm{rank}}Q_{a}=n-k+1-\dim K.$
Write
$r^{(i)}\equiv t^{i}\mod a,\quad\deg r^{(i)}<\deg a$
and
$b^{(i)}=\left\{\begin{array}[]{cl}t^{i}&\text{if }i<k,\\ t^{i}-r^{(i)}&\text{otherwise},\end{array}\right.\quad i=0,1,\dots,n.$
Clearly, for each polynomial $f\in{\mathcal{G}}(n)$ of form
$f=\sum_{i=0}^{n}f_{i}t^{i}$
one has
$f=\sum_{i=0}^{n}c_{i}b^{(i)}\quad\text{ for some }c_{0},c_{1},\dots,$
moreover
(7) $a\mid f\quad\Leftrightarrow\quad c_{0}=\dots=c_{k-1}=0,\text{ and }c_{k}=f_{k},\dots,c_{n}=f_{n}.$
Indeed, if $a\mid f$, then
$f=f-(f\bmod a)=\sum_{i=0}^{n}f_{i}(t^{i}-r^{(i)})=\sum_{i=k}^{n}f_{i}b^{(i)},$
as $t^{i}=r^{(i)}$ for $0\leq i<k$. Hence, $b^{(k)},\dots,b^{(n)}$ is a basis of the vector space of polynomials of form $ah$, $h\in{\mathcal{G}}(n-k+1)$.
Now assume, that $g\in K$ and write
$ag=\sum_{i=0}^{n}d_{i}t^{i}.$
By (7), we can also write
(8) $ag=\sum_{i=k}^{n}d_{i}b^{(i)}.$
As $g\in K$, we have by (5) that for $\ell=k,\dots,n$
$0$ $=B(ag,b^{(\ell)})=B\left(\sum_{i=0}^{n}d_{i}t^{i},t^{\ell}-r^{(\ell)}\right)$
(9) $=B\left(\sum_{i=0}^{n}d_{i}t^{i},t^{\ell}\right)-B\left(\sum_{i=0}^{n}d_{i}t^{i},r^{(\ell)}\right)$
$=\frac{d_{\ell+1}+d_{\ell-1}}{2}-B\left(\sum_{i=0}^{n}d_{i}t^{i},r^{(\ell)}\right).$
As $\deg r^{(\ell)}<k$, we have $B(t^{i},r^{(\ell)})=0$ for $i>k$. The (8) and (4.3) yield that
$d_{\ell+1}+d_{\ell-1}=2B\left(\sum_{i=0}^{n}d_{i}t^{i},r^{(\ell)}\right)=2d_{k}B\left(r^{(k)},r^{(\ell)}\right),\quad\text{for }k\leq\ell<n.$
That is, $d_{k}$ and $d_{k+1}$ determine all the values of $d_{k+2},\dots,d_{n}$. Also, as
$\sum_{i=0}^{k-1}d_{i}t^{i}=\sum_{i=k}^{n}d_{i}(t^{i}\bmod a),$
the values $d_{0},\dots,d_{k-1}$ are also uniquely determined. Thus $\dim K=2$. Then the result follows from (6). ∎
In the next lemma we investigate the rank of the quadratic form $h\mapsto S(ah)-S(bh)$ with $h\in{\mathcal{G}}(n-k)$.
Lemma 8.
Let $a,b\in{\mathcal{M}}(k)$ be polynomials such that
(10) $a^{*}\cdot a\neq b^{*}\cdot b.$
Then, the quadratic form $h\mapsto S(ah)-S(bh)$ with $h\in{\mathcal{G}}(n-k)$ has rank at least n2k1n-2k-1.
Proof.
Recall, that the symmetric bilinear form $B$ defined by (5) satisfies $S(h)=B(h,h)$.
First consider the symmetric bilinear form $B_{a}$ defined by
$B_{a}(h,g)=B(ah,ag),$
with $h,g\in{\mathcal{G}}(n-k)$. Then
$B_{a}(t^{i},t^{j})=B(a\cdot t^{i},a\cdot t^{j})=\frac{1}{2}\left(S^{(|i-j-1|)}(a)+S^{(|j-i-1|)}(a)\right).$
Similarly, if $B_{b}$ is the symmetric bilinear form defined by $B_{b}(h,g)=B(bh,bg)$, $h,g\in{\mathcal{G}}(n-k)$, we also have
$B_{b}(t^{i},t^{j})=B(b\cdot t^{i},b\cdot t^{j})=\frac{1}{2}\left(S^{(|i-j-1|)}(b)+S^{(|j-i-1|)}(b)\right).$
Whence, the symmetric bilinear form $B_{a,b}$ corresponding to S(ah)S(bh)S(ah)-S(bh) satisfies
(11) $B_{a,b}(t^{i},t^{j})=\frac{1}{2}\left(S^{(|i-j-1|)}(a)+S^{(|j-i-1|)}(a)-S^{(|i-j-1|)}(b)-S^{(|j-i-1|)}(b)\right).$
Let $M\in\mathbb{F}_{q}^{(n-k+1)\times(n-k+1)}$ be the symmetric matrix corresponding to $B_{a,b}$ with respect to the basis $1,t,t^{2},\dots,t^{n-k}$. We show, that it has rank at least n2k1n-2k-1. To do so let $\ell$ be the smallest integer such that $t^{\ell}\nmid a^{*}\cdot a-b^{*}\cdot b$. By (10), $\ell$ is well-defined. Also, as $a^{*}\cdot a-b^{*}\cdot b$ is symmetric, $\ell\leq k$. Then, by the definition of $\ell$ and by Lemma 5 we have
(12) $S^{(\nu)}(a)=S^{(\nu)}(b)\quad\text{for }k-\ell<\nu\leq k\quad\text{and}\quad S^{(k-\ell)}(a)\neq S^{(k-\ell)}(b).$
By definition of $S^{(\nu)}$, we also have, that
(13) $S^{(\nu)}(a)=S^{(\nu)}(b)=0\quad\text{for }\nu>k.$
Then the submatrix
$(M_{i,j})\quad\text{with }k-\ell+2\leq i\leq n-k+1,\quad 1\leq j\leq n-2k+\ell$
of $M$ is an upper triangular matrix with nonzero elements in the diagonal by (12) and (11). Indeed, for $0\leq\nu,\mu<n-2k+\ell$ with $\nu>\mu$, we have
$M_{k-\ell+2+\nu,1+\mu}=\frac{1}{2}\Big{(}$ $S^{(|k-\ell+\nu-\mu|)}(a)+S^{(|\mu-\nu-k+\ell-2|)}(a)$
$\quad-S^{(|k-\ell+\nu-\mu|)}(b)-S^{(|\mu-\nu-k+\ell-2|)}(b)\Big{)}=0.$
by (13) as $|k-\ell+\nu-\mu|>k$-$\ell$ and $|\mu-\nu-k+\ell-2|>k-\ell+2$. Moreover, the values in the diagonal are
$M_{k-\ell+2+\nu,1+\nu}$
$=\frac{1}{2}\left(S^{(|k-\ell|)}(a)+S^{(|-k+\ell-2|)}(a)-S^{(|k-\ell|)}(b)-S^{(|-k+\ell-2|)}(b)\right)$
$=\frac{1}{2}\left(S^{(k-\ell)}(a)-S^{(k-\ell)}(b)\right)\neq 0$
by (12) and (13) as $|-k+\ell-2|>k$-$\ell$.
Thus it has rank $n-2k+\ell-1$ and so
$\mathop{\mathrm{rank}}M\geq n-2k+\ell-1\geq n-2k-1.$
which proves the result. ∎
4.4. Vaughan’s identity for function fields
In this section we revise Vaughan’s identity for function fields. To do that, let us define the von Mangoldt function $\Lambda$ in $\mathbb{F}_{q}[t]$ by
$\Lambda(f)=\left\{\begin{array}[]{cl}\deg P&\text{if $f=P^{r}$ for some irreducible polynomial $P$;}\\ 0&\text{otherwise.}\end{array}\right$.
Then we have the following result.
Lemma 9.
Let $\Psi:\mathbb{F}_{q}[t]\rightarrow\mathbb{C}$ with $|\Psi(f)|\leq 1$. If $1\leq u,v\leq n$ two integers with $u+v<n$, then
$\sum_{f\in{\mathcal{M}}(n)}\Lambda(f)\Psi(f)\ll n\Sigma_{1}+n^{5/2}q^{n-(u+v)/2}\Sigma_{2}^{1/2}$
with
$\Sigma_{1}$ $=\sum_{|g|\leq q^{u+v}}\left|\sum_{|h|=q^{n}/|g|}\Psi(gh)\right|$
$\Sigma_{2}$ $=\max_{v\leq i\leq n-u}\max_{|g_{1}|=q^{n-i}}\sum_{|g_{2}|=q^{n-i}}\left|\sum_{|h|=q^{i}}\Psi(hg_{1})\overline{\Psi(hg_{2})}\right|,$
where the summations are taken over monic polynomials and the implied constant is absolute.
Proof.
Let $\mu$ be the Möbius function on $\mathbb{F}_{q}[t]$ defined by
[…]
Then, one can write
$\Lambda(f)=\sum_{a\mid f}\mu(a)\deg(f/a),$
where the sum is taken over only monic polynomials. Then
$\Lambda(f)=\sum_{\begin{subarray}{c}a\mid f\\ |a|\leq q^{u}\end{subarray}}\mu(a)\deg(f/a)-\underset{\begin{subarray}{c}ab\mid f\\ |a|\leq q^{u},|b|\leq q^{v}\end{subarray}}{\sum\ \sum}\mu(b)\Lambda(b)+\underset{\begin{subarray}{c}ab\mid f\\ |a|>q^{u},|b|>q^{v}\end{subarray}}{\sum\ \sum}\mu(a)\Lambda(b).$
It yields
$\sum_{f\in{\mathcal{M}}(n)}\Lambda(f)\Psi(f)=S_{1}-S_{2}+S_{3},$
where
$S_{1}$ $=\underset{\begin{subarray}{c}|ab|=q^{n}\\ |a|\leq q^{u}\end{subarray}}{\sum\ \sum}\mu(a)\deg(b)\Psi(ab),$
$S_{2}$ $=\underset{\begin{subarray}{c}|abc|=q^{n}\\ |a|\leq q^{u},|b|\leq q^{v}\end{subarray}}{\sum\ \sum\ \sum}\mu(a)\Lambda(b)\Psi(abc),$
$S_{3}$ $=\underset{\begin{subarray}{c}|abc|=q^{n}\\ |a|>q^{u},|b|>q^{v}\end{subarray}}{\sum\ \sum\sum}\mu(a)\Lambda(b)\Psi(abc).$
For $S_{1}$, we have
$S_{1}=\sum_{|a|\leq q^{u}}\mu(a)(n-\deg(a))\sum_{\begin{subarray}{c}|b|=q^{n}/|a|\end{subarray}}\Psi(ab)\ll n\sum_{|a|\leq q^{u}}\left|\sum_{\begin{subarray}{c}|b|=q^{n}/|b|\end{subarray}}\Psi(ab)\right|.$
For $S_{2}$, observe, that
$\sum_{b\mid z}\Lambda(b)=\deg(z).$
Then, we have
$\begin{split}S_{2}&=\sum_{r=n-u-v}^{n}\left(\underset{\begin{subarray}{c}|a|\leq q^{u}\,|b|\leq q^{v}\\ |ab|=q^{n-r}\end{subarray}}{\sum\ \sum}\mu(a)\Lambda(b)\right)\sum_{|c|=q^{r}}\Psi(abc)\\ &\ll(u+v)\sum_{|z|\leq q^{u+v}}\left|\sum_{|c|=q^{n}/|z|}\Psi(zc)\right|.\end{split}$
For $S_{3}$, observe, that
$\sum_{a\mid c}\mu(a)=\left\{\begin{array}[]{cl}1&\text{if $c=1$};\\ 0&\text{otherwise}.\end{array}\right$.
Then, we have
$S_{3}=\sum_{q^{v}\leq|b|\leq q^{n-u}}\Lambda(b)\sum_{|c|=q^{n}/|b|}\left(\sum_{\begin{subarray}{c}a\mid c\\ |a|>q^{u}\end{subarray}}\mu(a)\right)\Psi(bc).$
For $c\in\mathbb{F}_{q}[t]$, write
$\beta_{c}=\sum_{\begin{subarray}{c}a\mid c\\ |a|>q^{u}\end{subarray}}\mu(a).$
Then, by the Cauchy-Schwarz inequality,
$\left|S_{3}\right|^{2}\leq\left(\sum_{q^{v}\leq|b|\leq q^{n-u}}\Lambda(b)^{2}\right)\sum_{q^{v}\leq|b|\leq q^{n-u}}\left|\sum_{|c|=q^{n}/|b|}\beta_{c}\Psi(bc)\right|^{2}.$
For the first term, we have
$\sum_{q^{v}\leq|b|\leq q^{n-u}}\Lambda(b)^{2}\leq n\sum_{|b|\leq q^{n-u}}\Lambda(b)\leq nq^{n-u}.$
For the second term,
$\sum_{q^{v}\leq|b|\leq q^{n-u}}\left|\sum_{|c|=q^{n}/|b|}\beta_{c}\Psi(bc)\right|^{2}$
$=\sum_{v\leq i\leq n-u}\sum_{|c_{1}|=q^{n-i}}\sum_{|c_{2}|=q^{n-i}}\beta_{c_{1}}\overline{\beta_{c_{2}}}\sum_{|b|=q^{i}}\Psi(bc_{1})\overline{\Psi(bc_{2})}$
$\ll\sum_{v\leq i\leq n-u}\sum_{|c_{1}|=q^{n-i}}|\beta_{c_{1}}|^{2}\sum_{|c_{2}|=q^{n-i}}\left|\sum_{|b|=q^{i}}\Psi(bc_{1})\overline{\Psi(bc_{2})}\right|$
$\ll\sum_{v\leq i\leq n-u}\left(\sum_{|c_{1}|=q^{n-i}}|\beta_{c_{1}}|^{2}\right)\max_{|c_{1}|=q^{n-i}}\sum_{|c_{2}|=q^{n-i}}\left|\sum_{|b|=q^{i}}\Psi(bc_{1})\overline{\Psi(bc_{2})}\right|$
by $|\beta_{c_{1}}\overline{\beta_{c_{2}}}|\leq|\beta_{c_{1}}|^{2}/2+|\beta_{c_{2}}|^{2}/2$ and by the symmetry.
By Lemma 3,
$\sum_{|c_{1}|=q^{n-i}}|\beta_{c_{1}}|^{2}\ll\sum_{|c_{1}|=q^{n-i}}|\tau(c_{1})|^{2}\ll n^{3}q^{n-i}.$
Thus
$|S_{3}|^{2}\ll n^{5}q^{2n-u-v}\max_{v\leq i\leq n-u}\max_{|c_{1}|=q^{n-i}}\sum_{|c_{2}|=q^{n-i}}\left|\sum_{|b|=q^{i}}\Psi(bc_{1})\overline{\Psi(bc_{2})}\right|.$
5. Proof of Theorem 1
Let $\gamma\in\mathbb{F}_{q}$. By the orthogonality of additive characters, we have
${\mathcal{R}}(\gamma)=\frac{1}{q}\sum_{\psi}\sum_{f\in{\mathcal{P}}(n)}\psi(R(f)-\gamma),$
where the first summation is taken over all additive characters of $\mathbb{F}_{q}$. For the trivial additive character $\psi_{0}$, we have
$\frac{1}{q}\sum_{f\in{\mathcal{P}}(n)}\psi_{0}(R(f)-\gamma)=\frac{\#{\mathcal{P}}(n)}{q}$
which gives the main term.
The number of irreducible polynomials $f$ of degree $\deg f<n$ with $\deg f\mid n$ is $O(q^{n/2})$ by (1), thus for non-trivial additive characters $\psi\neq\psi_{0}$, we have
$\begin{split}\sum_{f\in{\mathcal{P}}(n)}\psi(R(f))&=\frac{1}{n}\sum_{f\in{\mathcal{P}}(n)}\Lambda(f)\psi(R(f))\\ &=\frac{1}{n}\sum_{f\in{\mathcal{M}}(n)}\Lambda(f)\psi(R(f))+O(q^{n/2}).\end{split}$
Then, by Lemma 9,
(14) $\sum_{f\in{\mathcal{M}}(n)}\Lambda(f)\psi(R(f))\ll n\Sigma_{1}+n^{5/2}q^{n-(u+v)/2}\Sigma_{2}^{1/2},$
where
$\Sigma_{1}=\sum_{|g|\leq q^{u+v}}\left|\sum_{|h|=q^{n}/|g|}\psi(R(gh))\right|$
and
$\Sigma_{2}=\max_{v\leq i\leq n-u}\max_{|g_{1}|=q^{n-i}}\sum_{|g_{2}|=q^{n-i}}\left|\sum_{|h|=q^{i}}\psi(R(hg_{1}))\overline{\psi(R(hg_{2}))}\right|$
with some $u,v$ to be chosen later.
In order to estimate $\Sigma_{1}$, observe, that by Lemma 7, the quadratic form $h\mapsto S(gh)$ with $\deg h\leq i$ has rank at least $n-\deg g-1$. If we consider it only on monic polynomials $h$, then its rank reduced by at most one and in this case $R(gh)=S(gh)+L_{g}(h)$ by (2) with some linear polynomial $L_{g}$ in terms of the coefficients of $h$. Then, by Lemma 4,
(15) $\Sigma_{1}$ $\ll\sum_{|g|\leq q^{u+v}}q^{n-\deg g-(n-\deg g-2)/2}$
$=\sum_{|g|\leq q^{u+v}}q^{(n-\deg g)/2+1}$
$\ll q^{(n+2)/2}\sum_{i\leq u+v}q^{i/2}$
$\ll q^{(n+u+v+2)/2}.$
In order to estimate $\Sigma_{2}$, let us fix $v\leq i\leq n-u$ and consider
$I(g_{1},g_{1})=\sum_{|h|=q^{i}}\psi(R(g_{1}h)-R(g_{2}h)).$
with monic polynomials $g_{1},g_{2}$ of degree nin-i. For given $g_{1}$, let
${\mathcal{B}}(g_{1})=\{g\in{\mathcal{M}}(n-i):g^{*}g=g_{1}^{*}g_{1}\}.$
If $g_{1}\not\in{\mathcal{B}}(g_{1})$, we have by Lemma 8, that the rank of the quadratic form $h\mapsto S(g_{1}h)-S(g_{2}h)$ with $\deg h\leq i$ is at least n2(ni)1=2in1n-2(n-i)-1=2i-n-1. If we consider it only on monic polynomials $h\in{\mathcal{M}}(i)$, then its rank is reduced by at most one and in this case $R(g_{1}h)-R(g_{2}h)=S(g_{1}h)-S(g_{2}h)+L_{g_{1},g_{2}}(h)$ with some linear polynomial $L_{g_{1},g_{2}}$ in the coefficients of $h$. Thus, in this case by Lemma 4 we get
$I(g_{1},g_{2})\ll q^{i-(2i-n-2)/2}=q^{n/2+1}.$
If $g_{2}\in{\mathcal{B}}(g_{1})$, we use the trivial bound
$I(g_{1},g_{2})\ll q^{i}.$
By Lemma 6 we have
${\mathcal{B}}(g_{1})\ll\min\left\{2^{n},q^{(2+\varepsilon)\frac{n}{\log_{q}n}}\right\}.$
By choosing $\varepsilon=1/2$, we get
${\mathcal{B}}(g_{1})\ll q^{n/14},$
where the implied constant is absolute. Then
(16) $\Sigma_{2}\ll\max_{v\leq i\leq n-u}\left(q^{n/14}q^{i}+q^{n-i}q^{n/2+1}\right)\ll q^{15n/14-u}+q^{3n/2-v+1}.$
Combining (14), (15) and (16), we get
$\sum_{f\in{\mathcal{M}}(n)}\Lambda(f)\psi(R(f))$ $\ll nq^{(n+u+v+2)/2}$
$\quad+n^{5/2}q^{n-(u+v)/2}\left(q^{15n/28-u/2}+q^{3n/4-v/2+1/2}\right)$
$\ll nq^{(n+u+v+2)/2}$
$\quad+n^{5/2}\left(q^{43n/28-u-v/2}+q^{7n/4-u/2-v+1/2}\right).$
Choosing
$u=\frac{3}{14}n\quad\text{and}\quad v=\frac{10}{14}n,$
we get
$\sum_{f\in{\mathcal{M}}(n)}\Lambda(f)\psi(R(f))$ $\ll n^{5/2}q^{27n/28+1}.$
Acknowledgment
The author wishes to thank Arne Winterhof for his valuable comments.
The author was was partially supported by NRDI (National Research Development and Innovation Office, Hungary) grant FK 142960 and by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.